# The Analysis of Willmore Surfaces and its Generalizations in Higher Dimensions

### v1

@@ IV.3 Local injectivity and integral densities
And since $\nabla\tilde{f}_{i}\rightarrow M$ in $L^{p}(B^{n})$ for any $p\in[1,\infty)$, we have
$\left\|\sqrt{\det\big(\nabla\tilde{f}_{i}^{T}\nabla\tilde{f}_{i}\big)}-\sqrt{\det\big(M^{T}M\big)}\right\|_{L^{1}(B^{n})}\xrightarrow[i\rightarrow\infty]{}0.$ (IV.39)
By Sobolev embedding we have $\tilde{f}_{i}\rightarrow\tilde{f}$ in $C^{0}(\overline{B^{n}})$  [^6] In general, $f_{r}$ defined in (IV.32) does not has a $C^{0}$ limit as $r\to 0$ under the same assumptions. For instance, we may take $\bar{f}(x_{1},x_{2})\coloneqq\begin{cases}0&\text{if }(x_{1},x_{2})=0,\\[5.69054pt] \Big(x_{1},x_{2},100^{-1}\,x_{1}\cos\big(\log(1-\log(x_{1}^{2}+x_{2}^{2}))\big)\Big)&\text{otherwise}.\end{cases}$ Then $\bar{f}\in W^{2,2}_{\textup{imm}}(D^{2},\mathbb{R}^{3})$ hence $\nabla\bar{f}\in VMO(D^{2})$, but for any $\omega\in\{(x_{1},x_{2})\in D^{2}:x_{1}\neq 0\}$, $\bar{f}_{r}(\omega)\coloneqq r^{-1}\bar{f}(r\omega)$ diverges as $r\rightarrow 0$. Moreover, we compute the contingent cone [8, Definition 4.1.1] and Clarke tangent cone [8, Definition 4.1.5] of $\bar{f}(D^{2})$ at 0: $T_{\bar{f}(D^{2})}(0)=\big\{(x_{1},x_{2},x_{3})\in\mathbb{R}^{3}:|x_{3}|\leq|x_{1}|\big\},\qquad C_{\bar{f}(D^{2})}(0)=\{0\}\times\mathbb{R}\times\{0\}.$ Neither of the two tangent cones above is a $2$-plane in $\mathbb{R}^{3}$. However, if we assume $f\colon D^{2}\rightarrow\mathbb{R}^{3}$ is a $C^{1}$ embedding, then it holds that $T_{f(D^{2})}(0)=C_{f(D^{2})}(0)=\text{Im}(\nabla f(0)).$ , and since $\mathcal{L}^{n}\left(\tilde{f}^{-1}\big(\partial B_{C_{0}^{-1}}(0)\big)\right)=0$, by [55, Theorem 3.2.11] we obtain
$\mathcal{L}^{n}\left(\tilde{f}_{i}^{-1}\big(B_{C_{0}^{-1}}(0)\big)\right)\rightarrow\mathcal{L}^{n}\left(\tilde{f}^{-1}\big(B_{C_{0}^{-1}}(0)\big)\right).$ (IV.40)
@@
$\begin{aligned} \mathcal{L}^{n}\left(\tilde{f}^{-1}\big(B_{C_{0}^{-1}}(0)\big)\right)&=C_{0}^{-n}\mathcal{L}^{n}\left\{\omega\in\mathbb{R}^{n}:\langle M^{T}M\omega,\omega\rangle\leq 1\right\}.\end{aligned}$ (IV.41)
@@
$\begin{aligned} \mathcal{L}^{n}\left\{\omega\in\mathbb{R}^{n}:\langle M^{T}M\omega,\omega\rangle\leq 1\right\}&=\mathcal{L}^{n}\left\{\omega\in\mathbb{R}^{n}:\langle DQ^{-1}\omega,Q^{-1}\omega\rangle\leq 1\right\}\\[7.11317pt] &=\mathcal{L}^{n}\left\{v\in\mathbb{R}^{n}:\langle Dv,v\rangle\leq 1\right\}\\ &=\mathcal{L}^{n}\bigg\{(v_{1},\dots,v_{n})\in\mathbb{R}^{n}:\sum_{j=1}^{n}\lambda_{j}v_{j}^{2}\leq 1\bigg\}.\end{aligned}$ (IV.42)
@@
$\begin{aligned} \mathcal{L}^{n}\bigg\{(v_{1},\dots,v_{n})\in\mathbb{R}^{n}:\sum_{j=1}^{n}\lambda_{j}v_{j}^{2}\leq 1\bigg\}&=\prod_{j=1}^{n}\lambda_{j}^{-\frac{1}{2}}|B^{n}|=\det(M^{T}M)^{-\frac{1}{2}}|B^{n}|.\end{aligned}$ (IV.43)
Combining (IV.38)–(IV.43), we finally conclude
@@
Thus, the limit (IV.31) is proved. ∎
@@ IV.4 Approximation by smooth immersions
In this section, we prove that every weak immersion $\vec{\Phi}$ can always be approximated by smooth immersions $\vec{\Phi}_{k}$ and that the conformal structures induced by $\vec{\Phi}_{k}$ converge to the one of $\vec{\Phi}$. To do so, we first prove in Section IV.4.1 that the conformal charts obtained by Coulomb frames converges. Combining this with an argument from Uhlenbeck, we prove in Section IV.4.2 that the conformal structures converges as well. We summarize the approximations results that we obtain in Section IV.4.3.
@@ IV.4.1 Convergence of isothermal coordinates
$\partial_{i}\big(a^{ij}_{k}\,\partial_{j}u_{k}\big)=f_{k}\quad\text{in }U.$ (IV.44)
@@
Let $V\Subset U$ be a bounded open subset with smooth boundary. Let $\xi\in C^{\infty}_{c}(U)$ be a cut-off function such that $\xi\geq 0$ on $U$ and $\xi=1$ on $V$. Denote by $\left\langle,\right\rangle$ the duality pairing between $W^{-1,2}(U)$ and $W^{1,2}_{0}(U)$. Using $(u_{k}-u)\,\xi$ as a test function in the equation (IV.44), we obtain
$\begin{aligned} &\Lambda^{-1}\int_{V}|\nabla(u_{k}-u)|^{2}\\ &\leq\int_{U}a^{ij}_{k}\,\xi\,\partial_{j}(u_{k}-u)\,\partial_{i}(u_{k}-u)\\ &=-\left\langle f_{k},\xi(u_{k}-u)\right\rangle-\int_{U}\big(a^{ij}_{k}\,\partial_{j}u_{k}\,\partial_{i}\xi\,(u_{k}-u)+a^{ij}_{k}\,\xi\,\partial_{j}u\,\partial_{i}(u_{k}-u)\big).\end{aligned}$ (IV.45)
@@
$\begin{aligned} \left|\left\langle f_{k},\xi(u_{k}-u)\right\rangle\right|&\leq\left|\left\langle f_{k}-f,\xi(u_{k}-u)\right\rangle\right|+\left|\left\langle f,\xi(u_{k}-u)\right\rangle\right|\\[8.53581pt] &\leq\left\|f_{k}-f\right\|_{W^{-1,2}(U)}\left\|\xi(u_{k}-u)\right\|_{W^{1,2}(U)}+\left|\left\langle f,\xi(u_{k}-u)\right\rangle\right|\\[5.69054pt] &\xrightarrow[k\to\infty]{}0.\end{aligned}$ (IV.46)
@@
$\begin{aligned} \left|\int_{U}a^{ij}_{k}\,\partial_{j}u_{k}\,\partial_{i}\xi\,(u_{k}-u)\right|&\leq\left\|a^{ij}_{k}\right\|_{L^{\infty}(U)}\left\|\partial_{j}u_{k}\right\|_{L^{2}(U)}\left\|\partial_{i}\xi\,(u_{k}-u)\right\|_{L^{2}(U)}\xrightarrow[k\to\infty]{}0.\end{aligned}$ (IV.47)
@@
$\begin{aligned} &\left|\int_{U}a^{ij}_{k}\,\xi\,\partial_{j}u\,\partial_{i}(u_{k}-u)\right|&\\ &\leq\left|\int_{U}a^{ij}\,\xi\,\partial_{j}u\,\partial_{i}(u_{k}-u)\right|+\int_{U}\left|(a^{ij}_{k}-a^{ij})\,\xi\,\partial_{j}u\,\partial_{i}(u_{k}-u)\right|\\ &\leq\left|\int_{U}a^{ij}\,\xi\,\partial_{j}u\,\partial_{i}(u_{k}-u)\right|+\left\|(a^{ij}_{k}-a^{ij})\,\xi\,\partial_{j}u\right\|_{L^{2}(U)}\left\|\partial_{i}(u_{k}-u)\right\|_{L^{2}(U)}\\ &\xrightarrow[k\to\infty]{}0.\end{aligned}$ (IV.48)
Combining (LABEL:est-diff-W12)–(LABEL:3rd-est-Duk-diff) we finally conclude $u_{k}\rightarrow u$ in $W^{1,2}_{\textup{loc}}(U)$. ∎
@@
$\Lambda^{-1}|X|_{\mathbb{R}^{2}}\leq|d(\vec{\Phi}_{k})_{p}(X)|_{\mathbb{R}^{m}}\leq\Lambda|X|_{\mathbb{R}^{2}}.$ (IV.49)
@@
$\vec{f}_{1,k}\coloneqq\frac{\partial_{1}\vec{\Phi}_{k}}{|\partial_{1}\vec{\Phi}_{k}|},\qquad\vec{f}_{2,k}\coloneqq\frac{\partial_{2}\vec{\Phi}_{k}-(\partial_{2}\vec{\Phi}_{k}\cdot\vec{f}_{1,k})\vec{f}_{1,k}}{|\partial_{2}\vec{\Phi}_{k}-(\partial_{2}\vec{\Phi}_{k}\cdot\vec{f}_{1,k})\vec{f}_{1,k}|}.$ (IV.50)
By (IV.49) we have $|\partial_{1}\vec{\Phi}_{k}|,|\partial_{2}\vec{\Phi}_{k}|\in[\Lambda^{-1},\Lambda]$, and
$|\partial_{2}\vec{\Phi}_{k}-(\partial_{2}\vec{\Phi}_{k}\cdot\vec{f}_{1,k})\vec{f}_{1,k}|\in[\Lambda^{-1},\Lambda(1+\Lambda^{2})].$
@@
$\begin{array}[]{l}\displaystyle\inf_{\theta\in W_{0}^{1,2}(D^{2},{\mathbb{R}})}\int_{D^{2}}|\langle\vec{f}^{\,\theta}_{1,k},d\vec{f}^{\,\theta}_{2,k}\rangle|^{2}_{g_{k}}\ d\textup{vol}_{g_{k}}\\[14.22636pt] \displaystyle=\inf_{\theta\in W_{0}^{1,2}(D^{2},{\mathbb{R}})}\int_{D^{2}}\sum_{i,j=1}^{2}g^{ij}_{k}\,\langle\vec{f}^{\,\theta}_{1,k},\partial_{i}\vec{f}^{\,\theta}_{2,k}\rangle\,\langle\vec{f}^{\,\theta}_{1,k},\partial_{j}\vec{f}^{\,\theta}_{2,k}\rangle\,\sqrt{\det g_{k}}\ dx^{1}\wedge dx^{2}.\end{array}$ (IV.51)
@@
$\langle\vec{f}^{\,\theta}_{1,k},\partial_{i}\vec{f}^{\,\theta}_{2,k}\rangle=\partial_{i}\theta+\langle\vec{f}_{1,k},\partial_{i}\vec{f}_{2,k}\rangle.$ (IV.52)
@@
$\forall\phi\in W_{0}^{1,2}(D^{2},{\mathbb{R}}),\qquad\int_{D^{2}}d\phi\wedge*_{g_{k}}\left(d\theta_{k}+\langle\vec{f}_{1,k},d\vec{f}_{2,k}\rangle\right)=0.$ (IV.53)
In particular, taking $\phi=\theta_{k}$ in (IV.53), we get
@@
Thus the sequence $\{\theta_{k}\}$ is bounded in $W_{0}^{1,2}(D^{2})$ and each subsequence contains a further subsequence $\{\theta_{k_{l}}\}$ weakly converging to some $\theta^{\prime}\in W^{1,2}_{0}(D^{2})$. Then using (IV.53), Hölder’s inequality and dominated convergence theorem, we have
@@
Then by replacing the cutoff function $\varphi$ by $1$ on $D^{2}$, we use the same argument as in Lemma IV.16 to obtain $\theta_{k}\rightarrow\theta_{\infty}$ in $W^{1,2}(D^{2})$. Denoting $(\vec{e}_{1,k},\vec{e}_{2,k})\coloneqq(\vec{f}^{\,\theta_{k}}_{1,k},\vec{f}^{\,\theta_{k}}_{2,k})$, we have $\vec{e}_{i,k}\rightarrow\vec{e}_{i,\infty}$ as $k\rightarrow\infty$ for $i\in\{1,2\}$. We deduce from (IV.53) that the following equation holds in the sense of distribution:
@@
$\begin{dcases}e_{i,k}\coloneqq g^{jl}_{k}\,\langle\vec{e}_{i,k},\partial_{j}\vec{\Phi}_{k}\rangle\,\frac{\partial}{\partial x^{l}}\in L^{\infty}\cap W^{1,2}(D_{1/2}(0),TD_{1/2}(0)),\\[5.69054pt] e_{i,k}^{\ast}\coloneqq\langle\vec{e}_{i,k},\partial_{j}\vec{\Phi_{k}}\rangle\,dx^{j}=\langle\vec{e}_{i,k},d\vec{\Phi}_{k}\rangle\in L^{\infty}\cap W^{1,2}(D_{1/2}(0),T^{*}D_{1/2}(0)).\end{dcases}$ (IV.54)
@@
$d\varphi^{i}_{k}=e^{-\lambda_{k}}\,e_{i,k}^{*}.$ (IV.55)
Let $\varphi_{k}\coloneqq(\varphi^{1}_{k},\varphi^{2}_{k})\in W^{1,\infty}\cap W^{2,2}(D_{1/2}(0),{\mathbb{R}^{2}})$. From (IV.54) and (IV.55) we deduce that for a.e. $p\in D_{1/2}(0)$ and any $v\in T_{p}(D_{1/2}(0))$,
$e^{-\|\lambda_{k}\|_{C(\overline{D_{1/2}(0)})}}\,\Lambda^{-1}|v|\leq|d(\varphi_{k})_{p}(v)|=e^{-\lambda_{k}}|d\vec{\Phi}_{k}(v)|\leq e^{\|\lambda_{k}\|_{C(\overline{D_{1/2}(0)})}}\,\Lambda|v|.$ (IV.56)
Since $\lambda_{k}\rightarrow\lambda_{\infty}$, $e_{i,k}^{*}\rightarrow e^{*}_{i,\infty}$ in $W^{1,2}(D^{2})$ as $k\rightarrow\infty$, and $\{\lambda_{k}\}$ is bounded in $C(\overline{D_{1/2}(0)})$, by dominated convergence theorem and Poincaré inequality we have $\varphi_{k}\rightarrow\varphi_{\infty}$ in $W^{2,2}(D_{1/2}(0))$. Combining (IV.56) and Lemma IV.12, it follows that there exists an open neighborhood $U\subset D_{1/2}(0)$ of 0 independent of $k$ such that $\varphi_{k}$ is open and injective on $U$, and its inverse $\varphi_{k}^{-1}\in W^{1,\infty}\cap W^{2,2}(\varphi_{k}(U),\mathbb{R}^{2})$. We obtain that $\varphi_{k}$ is an isothermal chart of $\vec{\Phi}_{k}$ by the same proof of Theorem IV.4. ∎
IV.4.2 Convergence of conformal structure

### v3

@@ IV.3 Local injectivity and integral densities
Since $\nabla\tilde{f}_{i}\rightarrow M$ in $L^{p}(B^{n})$ for any $p\in[1,\infty)$, we have
$\left\|\sqrt{\det\big(\nabla\tilde{f}_{i}^{T}\nabla\tilde{f}_{i}\big)}-\sqrt{\det\big(M^{T}M\big)}\right\|_{L^{1}(B^{n})}\xrightarrow[i\rightarrow\infty]{}0.$ (IV.40)
By Sobolev embedding we have $\tilde{f}_{i}\rightarrow\tilde{f}$ in $C^{0}(\overline{B^{n}})$  [^8] In general, $f_{r}$ defined in (IV.33) does not have a $C^{0}$ limit as $r\to 0$ under the same assumptions. For instance, we may take $\bar{f}(x_{1},x_{2})\coloneqq\begin{cases}0&\text{if }(x_{1},x_{2})=0,\\[5.69054pt] \Big(x_{1},x_{2},100^{-1}\,x_{1}\cos\big(\log(1-\log(x_{1}^{2}+x_{2}^{2}))\big)\Big)&\text{otherwise}.\end{cases}$ Then $\bar{f}\in W^{2,2}_{\textup{imm}}(D^{2},\mathbb{R}^{3})$ hence $\nabla\bar{f}\in VMO(D^{2})$, but for any $\omega\in\{(x_{1},x_{2})\in D^{2}:x_{1}\neq 0\}$, $\bar{f}_{r}(\omega)\coloneqq r^{-1}\bar{f}(r\omega)$ diverges as $r\rightarrow 0$. Moreover, we compute the contingent cone [8, Definition 4.1.1] and Clarke tangent cone [8, Definition 4.1.5] of $\bar{f}(D^{2})$ at 0: $T_{\bar{f}(D^{2})}(0)=\big\{(x_{1},x_{2},x_{3})\in\mathbb{R}^{3}:|x_{3}|\leq|x_{1}|\big\},\qquad C_{\bar{f}(D^{2})}(0)=\{0\}\times\mathbb{R}\times\{0\}.$ Neither of the two tangent cones above is a $2$-plane in $\mathbb{R}^{3}$. However, if we assume $f\colon D^{2}\rightarrow\mathbb{R}^{3}$ is a $C^{1}$ embedding, then it holds that $T_{f(D^{2})}(0)=C_{f(D^{2})}(0)=\text{Im}(\nabla f(0)).$ . Since $\mathcal{L}^{n}\left(\tilde{f}^{-1}\big(\partial B_{C_{0}^{-1}}(0)\big)\right)=0$, by [55, Theorem 3.2.11] we obtain
$\mathcal{L}^{n}\left(\tilde{f}_{i}^{-1}\big(B_{C_{0}^{-1}}(0)\big)\right)\rightarrow\mathcal{L}^{n}\left(\tilde{f}^{-1}\big(B_{C_{0}^{-1}}(0)\big)\right).$ (IV.41)
@@
$\begin{aligned} \mathcal{L}^{n}\left(\tilde{f}^{-1}\big(B_{C_{0}^{-1}}(0)\big)\right)&=C_{0}^{-n}\mathcal{L}^{n}\left\{\omega\in\mathbb{R}^{n}:\langle M^{T}M\omega,\omega\rangle\leq 1\right\}.\end{aligned}$ (IV.42)
@@
$\begin{aligned} \mathcal{L}^{n}\left\{\omega\in\mathbb{R}^{n}:\langle M^{T}M\omega,\omega\rangle\leq 1\right\}&=\mathcal{L}^{n}\left\{\omega\in\mathbb{R}^{n}:\langle DQ^{-1}\omega,Q^{-1}\omega\rangle\leq 1\right\}\\[7.11317pt] &=\mathcal{L}^{n}\left\{v\in\mathbb{R}^{n}:\langle Dv,v\rangle\leq 1\right\}\\ &=\mathcal{L}^{n}\bigg\{(v_{1},\dots,v_{n})\in\mathbb{R}^{n}:\sum_{j=1}^{n}\lambda_{j}v_{j}^{2}\leq 1\bigg\}.\end{aligned}$ (IV.43)
@@
$\begin{aligned} \mathcal{L}^{n}\bigg\{(v_{1},\dots,v_{n})\in\mathbb{R}^{n}:\sum_{j=1}^{n}\lambda_{j}v_{j}^{2}\leq 1\bigg\}&=\prod_{j=1}^{n}\lambda_{j}^{-\frac{1}{2}}|B^{n}|=\det(M^{T}M)^{-\frac{1}{2}}|B^{n}|.\end{aligned}$ (IV.44)
Combining (IV.39)–(IV.44), we finally conclude
@@
Thus, the limit (IV.32) is proved. ∎
@@ IV.4 Approximation by smooth immersions
In this section, we prove that every weak immersion $\vec{\Phi}$ can be approximated by smooth immersions $\vec{\Phi}_{k}$ and that the conformal structures induced by $\vec{\Phi}_{k}$ converge to that of $\vec{\Phi}$. To do so, we first prove in Section IV.4.1 that the conformal charts obtained by Coulomb frames converge. Combining this with an argument from Uhlenbeck, we prove in Section IV.4.2 that the conformal structures converge as well. We summarize the approximation results that we obtain in Section IV.4.3.
@@ IV.4.1 Convergence of isothermal coordinates
$\partial_{i}\big(a^{ij}_{k}\,\partial_{j}u_{k}\big)=f_{k}\quad\text{in }U.$ (IV.45)
@@
Let $V\Subset U$ be a bounded open subset with smooth boundary. Let $\xi\in C^{\infty}_{c}(U)$ be a cut-off function such that $\xi\geq 0$ on $U$ and $\xi=1$ on $V$. Denote by $\left\langle,\right\rangle$ the duality pairing between $W^{-1,2}(U)$ and $W^{1,2}_{0}(U)$. Using $(u_{k}-u)\,\xi$ as a test function in the equation (IV.45), we obtain
$\begin{aligned} &\Lambda^{-1}\int_{V}|\nabla(u_{k}-u)|^{2}\leq\int_{U}a^{ij}_{k}\,\xi\,\partial_{j}(u_{k}-u)\,\partial_{i}(u_{k}-u)\\ &\leq-\left\langle f_{k},\xi(u_{k}-u)\right\rangle-\int_{U}\big(a^{ij}_{k}\,\partial_{j}u_{k}\,\partial_{i}\xi\,(u_{k}-u)+a^{ij}_{k}\,\xi\,\partial_{j}u\,\partial_{i}(u_{k}-u)\big).\end{aligned}$ (IV.46)
@@
$\begin{aligned} \left|\left\langle f_{k},\xi(u_{k}-u)\right\rangle\right|&\leq\left|\left\langle f_{k}-f,\xi(u_{k}-u)\right\rangle\right|+\left|\left\langle f,\xi(u_{k}-u)\right\rangle\right|\\[8.53581pt] &\leq\left\|f_{k}-f\right\|_{W^{-1,2}(U)}\left\|\xi(u_{k}-u)\right\|_{W^{1,2}(U)}+\left|\left\langle f,\xi(u_{k}-u)\right\rangle\right|\\[5.69054pt] &\xrightarrow[k\to\infty]{}0.\end{aligned}$ (IV.47)
@@
$\begin{aligned} \left|\int_{U}a^{ij}_{k}\,\partial_{j}u_{k}\,\partial_{i}\xi\,(u_{k}-u)\right|&\leq\left\|a^{ij}_{k}\right\|_{L^{\infty}(U)}\left\|\partial_{j}u_{k}\right\|_{L^{2}(U)}\left\|\partial_{i}\xi\,(u_{k}-u)\right\|_{L^{2}(U)}\xrightarrow[k\to\infty]{}0.\end{aligned}$ (IV.48)
@@
$\begin{aligned} \left|\int_{U}a^{ij}_{k}\,\xi\,\partial_{j}u\,\partial_{i}(u_{k}-u)\right|&\leq\left|\int_{U}a^{ij}\,\xi\,\partial_{j}u\,\partial_{i}(u_{k}-u)\right|+\int_{U}\left|(a^{ij}_{k}-a^{ij})\,\xi\,\partial_{j}u\,\partial_{i}(u_{k}-u)\right|\\ &\leq\left|\int_{U}a^{ij}\,\xi\,\partial_{j}u\,\partial_{i}(u_{k}-u)\right|+\left\|(a^{ij}_{k}-a^{ij})\,\xi\,\partial_{j}u\right\|_{L^{2}(U)}\left\|\partial_{i}(u_{k}-u)\right\|_{L^{2}(U)}\\ &\xrightarrow[k\to\infty]{}0.\end{aligned}$ (IV.49)
Combining (LABEL:est-diff-W12)–(IV.49) we finally conclude $u_{k}\rightarrow u$ in $W^{1,2}_{\textup{loc}}(U)$. ∎
@@
$\Lambda^{-1}|X|_{\mathbb{R}^{2}}\leq|d(\vec{\Phi}_{k})_{p}(X)|_{\mathbb{R}^{m}}\leq\Lambda|X|_{\mathbb{R}^{2}}.$ (IV.50)
@@
$\vec{f}_{1,k}\coloneqq\frac{\partial_{1}\vec{\Phi}_{k}}{|\partial_{1}\vec{\Phi}_{k}|},\qquad\vec{f}_{2,k}\coloneqq\frac{\partial_{2}\vec{\Phi}_{k}-(\partial_{2}\vec{\Phi}_{k}\cdot\vec{f}_{1,k})\vec{f}_{1,k}}{|\partial_{2}\vec{\Phi}_{k}-(\partial_{2}\vec{\Phi}_{k}\cdot\vec{f}_{1,k})\vec{f}_{1,k}|}.$ (IV.51)
By (IV.50) we have $|\partial_{1}\vec{\Phi}_{k}|,|\partial_{2}\vec{\Phi}_{k}|\in[\Lambda^{-1},\Lambda]$, and
$\left|\partial_{2}\vec{\Phi}_{k}-(\partial_{2}\vec{\Phi}_{k}\cdot\vec{f}_{1,k})\vec{f}_{1,k}\right|\in[\Lambda^{-1},\Lambda(1+\Lambda^{2})].$
@@
$\begin{array}[]{l}\displaystyle\inf_{\theta\in W_{0}^{1,2}(D^{2},{\mathbb{R}})}\int_{D^{2}}|\langle\vec{f}^{\,\theta}_{1,k},d\vec{f}^{\,\theta}_{2,k}\rangle|^{2}_{g_{k}}\ d\textup{vol}_{g_{k}}\\[14.22636pt] \displaystyle=\inf_{\theta\in W_{0}^{1,2}(D^{2},{\mathbb{R}})}\int_{D^{2}}\sum_{i,j=1}^{2}g^{ij}_{k}\,\langle\vec{f}^{\,\theta}_{1,k},\partial_{i}\vec{f}^{\,\theta}_{2,k}\rangle\,\langle\vec{f}^{\,\theta}_{1,k},\partial_{j}\vec{f}^{\,\theta}_{2,k}\rangle\,\sqrt{\det g_{k}}\ dx^{1}\wedge dx^{2}.\end{array}$ (IV.52)
@@
$\langle\vec{f}^{\,\theta}_{1,k},\partial_{i}\vec{f}^{\,\theta}_{2,k}\rangle=\partial_{i}\theta+\langle\vec{f}_{1,k},\partial_{i}\vec{f}_{2,k}\rangle.$ (IV.53)
@@
$\forall\phi\in W_{0}^{1,2}(D^{2},{\mathbb{R}}),\qquad\int_{D^{2}}d\phi\wedge*_{g_{k}}\left(d\theta_{k}+\langle\vec{f}_{1,k},d\vec{f}_{2,k}\rangle\right)=0.$ (IV.54)
In particular, taking $\phi=\theta_{k}$ in (IV.54), we get
@@
Thus the sequence $\{\theta_{k}\}$ is bounded in $W_{0}^{1,2}(D^{2})$ and each subsequence contains a further subsequence $\{\theta_{k_{l}}\}$ weakly converging to some $\theta^{\prime}\in W^{1,2}_{0}(D^{2})$. Then using (IV.54), Hölder’s inequality and dominated convergence theorem, we have
@@
Then by replacing the cutoff function $\varphi$ by $1$ on $D^{2}$, we use the same argument as in Lemma IV.16 to obtain $\theta_{k}\rightarrow\theta_{\infty}$ in $W^{1,2}(D^{2})$. Denoting $(\vec{e}_{1,k},\vec{e}_{2,k})\coloneqq(\vec{f}^{\,\theta_{k}}_{1,k},\vec{f}^{\,\theta_{k}}_{2,k})$, we have $\vec{e}_{i,k}\rightarrow\vec{e}_{i,\infty}$ as $k\rightarrow\infty$ for $i\in\{1,2\}$. We deduce from (IV.54) that the following equation holds in the sense of distribution:
@@
$\begin{dcases}e_{i,k}\coloneqq g^{jl}_{k}\,\langle\vec{e}_{i,k},\partial_{j}\vec{\Phi}_{k}\rangle\,\frac{\partial}{\partial x^{l}}\in L^{\infty}\cap W^{1,2}(D_{1/2}(0),TD_{1/2}(0)),\\[5.69054pt] e_{i,k}^{\ast}\coloneqq\langle\vec{e}_{i,k},\partial_{j}\vec{\Phi_{k}}\rangle\,dx^{j}=\langle\vec{e}_{i,k},d\vec{\Phi}_{k}\rangle\in L^{\infty}\cap W^{1,2}(D_{1/2}(0),T^{*}D_{1/2}(0)).\end{dcases}$ (IV.55)
@@
$d\varphi^{i}_{k}=e^{-\lambda_{k}}\,e_{i,k}^{*}.$ (IV.56)
Let $\varphi_{k}\coloneqq(\varphi^{1}_{k},\varphi^{2}_{k})\in W^{1,\infty}\cap W^{2,2}(D_{1/2}(0),{\mathbb{R}^{2}})$. From (IV.55) and (IV.56) we deduce that for a.e. $p\in D_{1/2}(0)$ and any $v\in T_{p}(D_{1/2}(0))$,
$e^{-\|\lambda_{k}\|_{C(\overline{D_{1/2}(0)})}}\,\Lambda^{-1}|v|\leq|d(\varphi_{k})_{p}(v)|=e^{-\lambda_{k}}|d\vec{\Phi}_{k}(v)|\leq e^{\|\lambda_{k}\|_{C(\overline{D_{1/2}(0)})}}\,\Lambda|v|.$ (IV.57)
Since $\lambda_{k}\rightarrow\lambda_{\infty}$, $e_{i,k}^{*}\rightarrow e^{*}_{i,\infty}$ in $W^{1,2}(D^{2})$ as $k\rightarrow\infty$, and $\{\lambda_{k}\}$ is bounded in $C(\overline{D_{1/2}(0)})$, by dominated convergence theorem and Poincaré inequality we have $\varphi_{k}\rightarrow\varphi_{\infty}$ in $W^{2,2}(D_{1/2}(0))$. Combining (IV.57) and Lemma IV.12, it follows that there exists an open neighborhood $U\subset D_{1/2}(0)$ of 0 independent of $k$ such that $\varphi_{k}$ is open and injective on $U$, and its inverse $\varphi_{k}^{-1}\in W^{1,\infty}\cap W^{2,2}(\varphi_{k}(U),\mathbb{R}^{2})$. We obtain that $\varphi_{k}$ is an isothermal chart of $\vec{\Phi}_{k}$ by the same proof of Theorem IV.4. ∎
IV.4.2 Convergence of conformal structures